OpenStudy (anonymous):

MEDAL FOR BEST ANSWER. A tiger shark swims 1000 meters in 5 mins downstream. Upstream, it would take the same shark 20 mins to swim the same distance. What's the rate of the tiger shark still in water? What's the rate off the current? (This is a Wind and Current problem.) (Word problem involving systems.)

OpenStudy (anonymous):

1000=5(x+y) 1000=20(x-y) do you see how i got those? distance=speed*time let x be the speed of the shark in still water and y be the speed of the current

OpenStudy (anonymous):

5 and 20?

OpenStudy (anonymous):

Don't you have to convert it into a decimal? Because isn't it by hours?

OpenStudy (anonymous):

And wouldn't it be 5(x-y)....

OpenStudy (anonymous):

Well if you want mph then yes make them decimals. 5 mins with the current, so x+y

OpenStudy (anonymous):

It's 5 minutes downstream...

OpenStudy (anonymous):

Yes, so the shark is being helped by the current. Speed of the current+speed of the shark

OpenStudy (anonymous):

So with the minutes made into hours it will be 1000=1/12(x+y) 1000=1/3(x-y) now distribute on the right sides and solve

OpenStudy (anonymous):

I got a decimal... Still trying it out...

OpenStudy (anonymous):

For x or y?

OpenStudy (anonymous):

208.333333333 for the shark.

OpenStudy (anonymous):

1000=1/12x+1/12y 1000=1/3x-1/3y solve one for y 1000=1/12x+1/12y 1000-1/12x=1/12y 12000-x=y plug that into other equation for y 1000=1/3x-1/3(12000-x) solve for x

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

1000=4000 I'm pretty sure that's wrong.

OpenStudy (anonymous):

1000=1/3x-1/3(12000-x) 1000=1/3x-4000+1/3x 5000=2/3x 7500=x

OpenStudy (anonymous):

><''

OpenStudy (anonymous):

I see